# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v4

@@ VII SIMULATION RESULTS
Figure 2: (a) Trajectories obtained using our proposed approach. (b) Trajectories obtained using the separating hyperplane approach [12].
VIII CONCLUSION
@@
We propose a hybrid feedback controller for safe autonomous robot navigation in three-dimensional environments with arbitrarily-shaped convex obstacles. These obstacles may have nonsmooth boundaries, large sizes, and can be placed arbitrarily, provided they meet certain mild disjointedness requirements, as per Assumption 8. The proposed hybrid controller guarantees global asymptotic stability of the target location in the obstacle-free workspace. The obstacle-avoidance component of the control law relies on the projection of the robot’s center onto the nearest obstacle, enabling applications in a priori unknown environments, as discussed in Section VI. The proposed hybrid feedback control law generates discontinuous vector fields when switching between modes. Incorporating a smoothing mechanism in our proposed hybrid feedback would be an interesting practical extension. Moreover, extending our approach to second-order robot dynamics navigating in three-dimensional environments with non-convex obstacles, would be an interesting future work.
IX APPENDIX
IX-A Proof of Lemma 1
@@
Since the target location at the origin $\mathbf{0}$ belongs to the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}^{\circ}$, there exists some distance between the target location and the $r_{a}-$dilated obstacle $\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$. In other words, since $\mathbf{0}\in\mathcal{W}_{r_{a}}^{\circ},$ there exists $\bar{\delta}>0$ such that $d(\mathbf{0},\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))=\bar{\delta}.$ Notice that, since the workspace satisfies Assumption 8, according to (18), the set $\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ belongs to the landing region associated with obstacle $\mathcal{O}_{i}$ i.e., $\left(\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\right)\subset\overline{\mathcal{R}}_{l}^{i}.$ According to (15), the location $\Pi(\mathbf{0},\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ does not belong to the set ¯$\overline{\mathcal{R}}_{l}^{i}$. As a result, one has $d(\mathbf{0},\mathcal{R}_{l}^{i})>\bar{\delta}.$ Hence, it is clear that $\mathcal{B}_{\bar{\delta}}(\mathbf{0})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{J}_{0}^{\mathcal{W}}=\emptyset,$ where $\mathcal{B}_{\bar{\delta}}(\mathbf{0})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\neq\emptyset.$ Hence, one can set $\bar{\epsilon}\in(0,d(\mathbf{0},\mathcal{R}_{l}^{i})-\bar{\delta}]$ to ensure that $\mathcal{ER}^{\mathbf{h}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\neq\emptyset$ for all $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, for any $\epsilon\in(0,\bar{\epsilon}]$, where the set $\mathcal{ER}^{\mathbf{h}}$ is defined in (22).
IX-B Proof of Lemma 2
@@
The flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, defined in (25) are by construction closed subsets of $\mathbb{R}^{3}\times\mathbb{R}^{3}\times\mathbb{R}^{3}\times\mathbb{R}\times\mathbb{R}.$ Hence, condition 1 in Lemma 2 is satisfied.
@@
IX-C Proof sketch for Lemma 3
Since the proposed hybrid closed-loop system (31) satisfies the hybrid basic conditions outlined in Lemma 2, we make use of [20, Proposition 6.10] to show the forward invariance of $\mathcal{K}$. Specifically, we verify the satisfaction of the following viability condition:
@@ IX-C Proof sketch for Lemma 3
which allows us to establish the completeness of the solution $\xi$ to the hybrid closed-loop system (31). In (35), $\mathbf{T}_{\mathcal{F}}(\xi)$ represents the tangent cone to the set $\mathcal{F}$ at $\xi$.
@@
According to [20, Proposition 6.10], since (35) holds for all $\xi\in\mathcal{F}\setminus\mathcal{J}$, there exists a maximal solution to (31) for each initial condition in $\mathcal{K}$. Finite escape time can only occur through flow, but it cannot occur for $\mathbf{x}$ in the set $\mathcal{F}_{1}^{\mathcal{W}}$, as this set is bounded by definition (23), nor for $\mathbf{x}$ in the set $\mathcal{F}_{0}^{\mathcal{W}}$, as this would make $\mathbf{x}^{\intercal}\mathbf{x}$ grow unbounded, contradicting the fact that $\frac{d}{dt}(\mathbf{x}^{\intercal}\mathbf{x})\leq 0$ according to the definition of $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)$. Therefore, all maximal solutions do not have finite escape times, and Claim (b) in [20, Proposition 6.10] does not hold.
According to (20) and (24), by construction, we have $\mathcal{F}_{m}\cup\mathcal{J}_{m}=\mathcal{K}$ for each $m\in\mathbb{M}$. Therefore, using (27), (30) and (31), it can be verified that $\mathbf{L}(\xi)\in\mathcal{F}\cup\mathcal{J}=\mathcal{K}$ for every $\xi\in\mathcal{K}$. As a result, Claim (c) in [20, Proposition 6.10] does not apply. Therefore, Claim (a) in [20, Proposition 6.10] holds. In other words, all maximal solutions to the hybrid closed-loop system (31) are complete, implying the forward invariance of the obstacle-free set $\mathcal{K}$.
IX-D Proof of Theorem 1
@@
Forward invariance and stability: The forward invariance of the obstacle-free set $\mathcal{K}$, for the hybrid closed-loop system (31), is immediate from Lemma 3. We next prove the stability of $\mathcal{A}$ using [20, Definition 7.1].

### v5

@@ VIII Simulation Results
Figure 12: (a) Robot trajectories obtained using our proposed hybrid feedback approach. (b) Robot trajectories obtained using the separating hyperplane approach [15].
Figure 13: Robot trajectories safely navigating around spherical obstacles and converging to the target location at the origin.
In the next simulation scenario, as shown in Fig. 13, we consider 2 three-dimensional spherical obstacles and apply the proposed hybrid controller (12) with modifications mentioned in Section VI, for a point robot initialized at $8$ different locations in the obstacle-free workspace. The safety distance $r_{s}$ is set to $0.15m$ and the parameter $\gamma=0.15m.$ The gains $\kappa_{s}$ and $\kappa_{r}$, used in (12a), are set to $0.5$. The sensing radius $R_{s}$, used in (8), is set to $2m$. The parameter $\epsilon$, used in (28), is set to $0.05m$. Similar to the previous simulations, it can be observed from Fig. 14 that the robot converges asymptotically to the target location without colliding with the obstacles. Since the obstacles are spheres, the distance between the robot and the target location monotonically decreases as the robot converges to the target location, as stated in Theorem 2 and as shown in Fig. 14b.
Figure 14: (a) Distance between the robot $\mathbf{x}$ and the obstacles $\mathcal{O}_{\mathcal{W}}$ versus time. (b) Distance between the robot $\mathbf{x}$ and the target location versus time.
Figure 15: (a) Robot trajectory for the hybrid closed-loop system (38) using the control input vector $\mathbf{u}(\xi)$. (b) Robot trajectory for the smoothed hybrid closed-loop system (50) using the control input vector $\mathbf{u}_{sm}(\xi,\tau)$.
We compare the robot trajectories generated by the hybrid closed-loop system (38) with the smoothed robot trajectories obtained from the modified hybrid closed-loop system (50). The environment contains two convex obstacles and the target location is positioned at the origin, as shown in Fig. 15. The robot, with a radius of $r=0.175m$, is initialized at the point $[5.5,3]^{\top}$. The safety distance $r_{s}$ is set to $0.025m$, and the parameter $\gamma$ is set to $0.3m$. The control gains $\kappa_{s}$ and $\kappa_{r}$ are set to 0.2 and 2, respectively. The sensing radius $R_{s}$ is set to $3m$, and the parameter $\epsilon$ is set to $1m$. The time parameters $t_{s}$ and $\tau_{s}$, used in (46), are set to $0.5$ seconds and $1$ second, respectively.
In Fig. 15, the red portions of the trajectories represent the motion of the robot in the move-to-target mode, whereas the blue portions correspond to the motion in the obstacle-avoidance mode. Figure 16 shows the time-evolution of the magnitude of the control input, while Fig. 17 depicts the direction of the control input represented by the angle between the control input vector and the $x$-axis. Given $\mathbf{p}=[p_{1},p_{2}]^{\top}\in\mathbb{R}^{2}$, the angle between $\mathbf{p}$ and the $x$-axis is computed as:
$\mathbf{atan2v}(\mathbf{p})=\mathbf{mod}(\mathbf{atan2}(p_{2}/p_{1})+2\pi,2\pi).$ (51)
In Fig. 16 and Fig. 17, the red trajectories represent the control input vector $\mathbf{u}$, while the blue trajectories correspond to $\mathbf{u}_{sm}$.
Generally, the magnitude and direction of the control input vector $\mathbf{u}$ changes instantaneously when switching between modes, as illustrated in Fig. 16 and Fig. 17, respectively. This can sometimes cause an abrupt change in the robot’s direction of motion when the control input $\mathbf{u}$ switches between modes, as shown in Fig. 15a. On the other hand, the magnitude and direction of the control input vector $\mathbf{u}_{sm}$, used in the modified hybrid closed-loop system (50), change continuously, as depicted in Fig. 16 and Fig. 17, respectively. Therefore, the control input vector $\mathbf{u}_{sm}$ generates smooth robot trajectories, as shown in Fig. 15b.
Figure 16: Magnitude of the control input vector versus time. The red trajectory represents $\|\mathbf{u}\|$, while the blue trajectory represents $\|\mathbf{u}_{sm}\|$.
Figure 17: Plot of the angle between the control input vectors and the positive $x$-axis versus time. The red trajectory represents $\mathbf{atan2v}(\mathbf{u})$, while the blue trajectory represents $\mathbf{atan2v}(\mathbf{u}_{sm})$.
IX Conclusion
@@
We propose a hybrid feedback controller for safe autonomous robot navigation in $n-$dimensional environments with arbitrarily-shaped convex obstacles. These obstacles may have nonsmooth boundaries and large sizes, and can be arbitrarily located, provided they meet certain mild disjointedness requirements, as per Assumption 9. The proposed hybrid controller guarantees global asymptotic stability of the target location in the obstacle-free workspace. The obstacle-avoidance component of the control law relies on the projection of the robot’s center onto the obstacle being avoided, enabling applications in a priori unknown environments, as discussed in Section VII. The proposed hybrid feedback control law generates discontinuous control inputs when switching between modes. A smoothing mechanism has been suggested to overcome this problem in practical applications. Extending our approach to robots, with second-order dynamics, navigating in $n-$dimensional environments with non-convex obstacles would be an interesting future work.
-A Proof of Lemma 1
First, we show that for any $R_{s}>0$, it holds that $\mathcal{R}_{a}^{i}\subset\mathcal{R}_{l}^{i}$. As shown in Fig. 18, let us partition $\mathcal{R}_{a}^{i}$ into two mutually exclusive sets as follows:
$\mathcal{R}_{a}^{i}=\mathcal{R}_{1}\cup\mathcal{R}_{2},$ (52)
where the sets $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are given by
$\mathcal{R}_{1}$ =$=\{\mathbf{x}\in\mathcal{R}_{a}^{i}|\eth_{i}^{\mathbf{x}}\cap\left(\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\cap(\mathcal{D}_{r_{a}}(\mathcal{L}_{s}(\mathbf{x},\mathbf{0})))^{\circ}\right)\neq\emptyset\},$
$\mathcal{R}_{2}$ $=\mathcal{R}_{a}^{i}\setminus\mathcal{R}_{1}.$
When $\mathbf{x}\in\mathcal{R}_{1}$, it is straightforward to notice that $\mathbf{x}\in\mathcal{R}_{l}^{i}$, where $\mathcal{R}_{l}^{i}$ is defined in (20). Therefore, we proceed to prove that $\mathbf{x}\in\mathcal{R}_{2}\implies\mathbf{x}\in\mathcal{R}_{l}^{i}$. Since $\mathcal{O}_{i}$ is a convex obstacle, when $\mathbf{x}\in\mathcal{R}_{2}$, one has $d(\mathcal{L}_{s}(\mathbf{x},\mathbf{0}),\eth_{i}^{\mathbf{x}})=d(\mathcal{L}_{s}(\mathbf{x},\mathbf{0}),\mathcal{O}_{i})=r_{a}$. Now, if we show that $\mathbf{x}^{\top}\mathbf{x}_{\pi}\geq 0$ for all $\mathbf{x}\in\mathcal{R}_{2}$, then, as per (20), one can conclude that $\mathbf{x}\in\mathcal{R}_{2}\implies\mathbf{x}\in\mathcal{R}_{l}^{i}$.
For $\mathbf{x}\in\mathcal{R}_{2}$, there exists $\mathbf{p}\in\eth_{i}^{\mathbf{x}}$ such that $\mathbf{p}\in\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})$. Therefore, it is true that $\mathbf{p}\in\mathcal{H}_{\leq}(\mathbf{x},\mathbf{x})\cap\partial\mathcal{O}_{i}$. Now, let us assume that $\mathbf{x}^{\top}\mathbf{x}_{\pi}<0$, which implies that $\Pi(\mathbf{x},\mathcal{O}_{i})\in\mathcal{H}_{>}(\mathbf{x},\mathbf{x})$. Since $\mathbf{p}\in\partial\mathcal{O}_{i}$ and $\Pi(\mathbf{x},\mathcal{O}_{i})\in\partial\mathcal{O}_{i}$, and obstacle $\mathcal{O}_{i}$ is convex, one has $\mathcal{L}_{s}(\mathbf{p},\Pi(\mathbf{x},\mathcal{O}_{i}))\subset\mathcal{O}_{i}$. Furthermore, since $\mathbf{p}\in\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\cap\mathcal{H}_{\leq}(\mathbf{x},\mathbf{x})$ and $\Pi(\mathbf{x},\mathcal{O}_{i})\in\mathcal{H}_{>}(\mathbf{x},\mathbf{x})$, it is true that $\mathcal{L}_{s}(\mathbf{p},\Pi(\mathbf{x},\mathcal{O}_{i}))\cap\left(\mathcal{B}_{\|\mathbf{x}_{\pi}\|}(\mathbf{x})\right)^{\circ}\neq\emptyset$. This implies that there exists $\mathbf{q}\in\mathcal{L}_{s}(\mathbf{p},\Pi(\mathbf{x},\mathcal{O}_{i}))\cap\mathcal{O}_{i}$ such that $\|\mathbf{x}-\mathbf{q}\|<\|\mathbf{x}_{\pi}\|$, which is a contradiction. Therefore, for $\mathbf{x}\in\mathcal{R}_{2}$, one can conclude that $\mathbf{x}^{\top}\mathbf{x}_{\pi}\geq 0$, and hence it is proved that for any $R_{s}>0$, $\mathcal{R}_{a}^{i}\subset\mathcal{R}_{l}^{i}$.
Next, we prove that if $R_{s}>l_{i}$, then $\mathcal{R}_{a}^{i}=\mathcal{R}_{l}^{i}$. Notice that the implication $\mathbf{x}\in\mathcal{R}_{a}^{i}\implies\mathbf{x}\in\mathcal{R}_{l}^{i}$ has been proved earlier for any $R_{s}>0$. Therefore, we focus on the backward implication and show that when $R_{s}>l_{i}$, $\mathbf{x}\in\mathcal{R}_{l}^{i}\implies\mathbf{x}\in\mathcal{R}_{a}^{i}$. For $\mathbf{x}\in\mathcal{R}_{l}^{i}$, there exists $\mathbf{p}\in\partial\mathcal{O}_{i}$ such that $d(\mathbf{p},\mathcal{L}_{s}(\mathbf{x},\mathbf{0}))\leq r_{a}$, $\mathcal{L}_{s}(\mathbf{x},\mathbf{p})\cap\mathcal{O}_{i}=\mathbf{p}$ and $\mathbf{x}^{\top}(\mathbf{x}-\mathbf{p})\geq 0$. Since $R_{s}>l_{i}$, it is straightforward to notice that $\mathbf{p}\in\eth_{i}^{\mathbf{x}}$. Therefore, according to (19), one can conclude that when $R_{s}>l_{i}$, $\mathbf{x}\in\mathcal{R}_{l}^{i}\implies\mathbf{x}\in\mathcal{R}_{a}^{i}$.
Figure 18: The representation of the sets $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$.
-B Proof of Lemma 2
@@
Since the target location at the origin $\mathbf{0}$ belongs to the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}^{\circ}$, there exists some distance between the target location and the $r_{a}-$dilated obstacle $\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$. In other words, since $\mathbf{0}\in\mathcal{W}_{r_{a}}^{\circ},$ there exists $\bar{\delta}>0$ such that $d(\mathbf{0},\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))=\bar{\delta}.$ Notice that, according to (24), the set $\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ belongs to the avoidance region associated with obstacle $\mathcal{O}_{i}$ i.e., $\left(\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\right)\subset\mathcal{R}_{a}^{i}$. In addition, as per Lemma 1, $\mathcal{R}_{a}^{i}\subset\mathcal{R}_{l}^{i}$, where $\mathcal{R}_{l}^{i}$ is defined in (20). Therefore, it is clear that $\left(\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\right)\subset\mathcal{R}_{l}^{i}$. According to (20), $\Pi(\mathbf{0},\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ does not belong to the set $\mathcal{R}_{l}^{i}$. As a result, one has $d(\mathbf{0},\mathcal{R}_{l}^{i})>\bar{\delta}.$ Hence, it is clear that $\mathcal{B}_{\bar{\delta}}(\mathbf{0})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{J}_{0}^{\mathcal{W}}=\emptyset,$ where $\mathcal{B}_{\bar{\delta}}(\mathbf{0})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\neq\emptyset.$ Hence, one can set $\bar{\epsilon}\in(0,d(\mathbf{0},\mathcal{R}_{l}^{i})-\bar{\delta}]$ to ensure that $\mathcal{ER}^{\mathbf{h}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\neq\emptyset$ for all $\mathbf{h}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, for any $\epsilon\in(0,\bar{\epsilon}]$, where the set $\mathcal{ER}^{\mathbf{h}}$ is defined in (28).
-C Proof of Lemma 3
@@
The flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, defined in (32) are by construction closed subsets of $\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}\times\mathbb{R}.$ Hence, condition 1 in Lemma 3 is satisfied.
@@
-D Proof of Lemma 4
First, we prove that the union of the flow and jump sets covers exactly the obstacle-free state space $\mathcal{K}$. For $m=0$, according to (24) and (25), by construction we have $\mathcal{F}_{0}^{\mathcal{W}}\cup\mathcal{J}_{0}^{\mathcal{W}}=\mathcal{W}_{r_{a}}.$ Similarly, for $m=1$, according to (27) and (30), by construction one has $\mathcal{F}_{1}^{\mathcal{W}}\cup\mathcal{J}_{1}^{\mathcal{W}}=\mathcal{W}_{r_{a}}$. Inspired by [22, Appendix 11], the satisfaction of the following equation:
$\mathcal{F}_{m}^{\mathcal{W}}\cup\mathcal{J}_{m}^{\mathcal{W}}=\mathcal{W}_{r_{a}},m\in\mathbb{M},$ (53)
along with (26), (31) and (32) implies $\mathcal{F}\cup\mathcal{J}=\mathcal{K}.$
Now, inspired by [22, Appendix 1], for the hybrid closed-loop system (38), with data $\mathcal{H}_{\mathcal{S}}=(\mathcal{F},\mathbf{F},\mathcal{J},\mathbf{J}),$ define $\mathbf{S}_{\mathcal{H}_{\mathcal{S}}}(\mathcal{K})$ as the set of all maximal solutions $\xi$ to $\mathcal{H}_{\mathcal{S}}$ with $\xi(0,0)\in\mathcal{K}$. Since $\mathcal{F}\cup\mathcal{J}=\mathcal{K}$, each $\xi\in\mathbf{S}_{\mathcal{H}_{\mathcal{S}}}(\mathcal{K})$ satisfies rge $\text{rge }\xi\subset\mathcal{K}.$ Additionally, if every maximal solution $\xi\in\mathbf{S}_{\mathcal{H}_{\mathcal{S}}}(\mathcal{K})$ is complete, then the set $\mathcal{K}$ will be forward invariant [29, Definition 3.13]. Since the hybrid closed-loop system (38) satisfies the hybrid basic conditions, as stated in Lemma 3, one can use [27, Proposition 6.10], to verify the following viability condition:
@@
which will allow us to establish the completeness of the solution $\xi$ to the hybrid closed-loop system (38). In (54), $\mathbf{T}_{\mathcal{F}}(\xi)$ represents the tangent cone to the set $\mathcal{F}$ at $\xi$.
Let $(\mathbf{x},\mathbf{h},\mathbf{a},m,s)\in\mathcal{F}\setminus\mathcal{J}$, which implies by virtue of (26), (31) and (32) that $(\mathbf{x},\mathbf{h},\mathbf{a},s)\in(\mathcal{F}_{m}^{\mathcal{W}}\setminus\mathcal{J}_{m}^{\mathcal{W}})\times\mathcal{W}_{r_{a}}\times\mathbb{S}^{n-1}\times\mathbb{R}_{\geq 0}$ for some $m\in\mathbb{M}$. For $\mathbf{x}\in(\mathcal{F}_{m}^{\mathcal{W}})^{\circ}\setminus\mathcal{J}_{m}^{\mathcal{W}}$ with $(\mathbf{h},\mathbf{a},m,s)\in\mathcal{W}_{r_{a}}\times\mathbb{S}^{n-1}\times\mathbb{M}\times\mathbb{R}_{\geq 0}$ the tangent cone $\mathbf{T}_{\mathcal{F}}(\xi)=\mathbb{R}^{n}\times\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})\times\mathcal{H}(\mathbf{0},\mathbf{a})\times\{0\}\times\mathbf{T}_{\mathbb{R}_{\geq 0}}(s),$ where the set $\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ is given by
$\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})=\begin{cases}\begin{matrix}[l]\mathbb{R}^{n},&\text{if }\mathbf{h}\in(\mathcal{W}_{r_{a}})^{\circ},\\ \mathcal{H}_{\geq}(\mathbf{0},(\mathbf{h}-\Pi(\mathbf{h},\mathcal{O}_{\mathcal{W}}))),&\text{if }\mathbf{h}\in\partial\mathcal{W}_{r_{a}},\end{matrix}\end{cases}$ (55)
where for $\mathbf{h}\in\partial\mathcal{W}_{r_{a}}$, the projection $\Pi(\mathbf{h},\mathcal{O}_{\mathcal{W}})$ is unique. For $s\in\mathbb{R}_{\geq 0}$, the set $\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ is defined as
$\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)=\begin{cases}\begin{matrix}[l]\mathbb{R},&\text{if }s\in\mathbb{R}_{>0},\\ \mathbb{R}_{\geq 0},&\text{if }s=0.\end{matrix}\end{cases}$ (56)
Since, according to (38), $\dot{\mathbf{h}}=\mathbf{0}$ and $\dot{s}=1$, we have $\dot{\mathbf{h}}\in\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ and $\dot{s}\in\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$, respectively, and (54) holds.
Next, we consider the case where $\xi\in\mathcal{F}\setminus\mathcal{J}$ with $\mathbf{x}\in\partial\mathcal{F}_{m}^{\mathcal{W}}\setminus\mathcal{J}_{m}^{\mathcal{W}}$, $m\in\mathbb{M}$. For $m=0$, according to (24) and (25), one has
$\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}}\subset\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}})\cap\mathcal{R}_{e},$ (57)
and for $\mathbf{x}\in\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}},$ the projection $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ is unique. Hence, for all $\xi\in\mathcal{F}_{0}\setminus\mathcal{J}_{0}$ with $\mathbf{x}\in\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}}$,
$\mathbf{T}_{\mathcal{F}}(\xi)=\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\times\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})\times\mathcal{H}(\mathbf{0},\mathbf{a})\times\{0\}\times\mathbf{T}_{\mathbb{R}_{\geq 0}}(s),$ (58)
where $\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ and $\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ are defined in (55) and (56), respectively, and $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$. Also, according to (12a), for $m=0$, one has $\mathbf{u}(\xi)=-\kappa_{s}\mathbf{x},\kappa_{s}>0$. According to (23) and (57), for $\xi\in\mathcal{F}_{0}\setminus\mathcal{J}_{0}$ with $\mathbf{x}\in\partial\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{J}_{0}^{\mathcal{W}}$, one can conclude that $\mathbf{u}(\xi)\in\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})$. Moreover, according to (38), it is clear that $\dot{\mathbf{h}}=\mathbf{0}\in\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$, $\dot{s}=1\in\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ and $\dot{\mathbf{a}}=\mathbf{0}\in\mathcal{H}(\mathbf{0},\mathbf{a})$. Therefore, the viability condition (54) holds for $m=0$.
For $m=1$, according to (27) and (30) one has
$\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}\subset\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}),$ (59)
and for $\mathbf{x}\in\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$, the projection $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ is unique and the set $\mathcal{B}_{\lVert\mathbf{x}_{\pi}\rVert}(\mathbf{x})$ intersects with $\partial\mathcal{O}_{\mathcal{W}}$ only at $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}}).$ Hence, for all $\xi\in\mathcal{F}_{1}\setminus\mathcal{J}_{1}$ with $\mathbf{x}\in\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$,
$\mathbf{T}_{\mathcal{F}}(\xi)=\mathcal{H}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\times\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})\times\mathcal{H}(\mathbf{0},\mathbf{a})\times\{0\}\times\mathbf{T}_{\mathbb{R}_{\geq 0}}(s),$ (60)
where $\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$ and $\mathbf{T}_{\mathbb{R}_{\geq 0}}(s)$ are defined in (55) and (56), respectively. Also, according to (12a), for $m=1$, one has $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a}),\kappa_{r}>0.$ From (16), it follows that $\eta(\mathbf{x})=1$ for $\mathbf{x}\in\partial\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$. As a result, according to (12a) and (13), the control vector is simplified to $\mathbf{u}(\xi)=\kappa_{r}{\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}}.$ Since for any two orthonormal vectors $\hat{\mathbf{h}},\mathbf{a}\in\mathbb{S}^{n-1}$, the matrix $\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})$ is positive semidefinite, one has $\mathbf{x}_{\pi}^{\intercal}\mathbf{P}(\hat{\mathbf{h}},\mathbf{a})\mathbf{x}_{\pi}\geq 0$. Therefore, for $\xi\in\mathcal{F}_{1}\setminus\mathcal{J}_{1}$ with $\mathbf{x}\in\mathcal{F}_{1}^{\mathcal{W}}\setminus\mathcal{J}_{1}^{\mathcal{W}}$, one has $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},1,s)^{\intercal}(\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}}))\geq 0$. Moreover, according to (38), it is clear that $\dot{\mathbf{h}}=\mathbf{0}\in\mathbf{T}_{\mathcal{W}_{r_{a}}}(\mathbf{h})$, $\dot{s}=1$ and $\dot{\mathbf{a}}=\mathbf{0}\in\mathcal{H}(\mathbf{0},\mathbf{a})$. Hence, the viability condition in (54) holds for $m=1.$
@@
Hence, according to [27, Proposition 6.10], since (54) holds for all $\xi\in\mathcal{F}\setminus\mathcal{J}$, there exists a nontrivial solution to $\mathcal{H}$ for each initial condition in $\mathcal{K}$. Finite escape time can only occur through flow. They can neither occur for $\mathbf{x}$ in the set $\mathcal{F}_{1}^{\mathcal{W}}$, as this set is bounded as per definition (30), nor for $\mathbf{x}$ in the set $\mathcal{F}_{0}^{\mathcal{W}}$ as this would make $\mathbf{x}^{\intercal}\mathbf{x}$ grow unbounded, and would contradict the fact that $\frac{d}{dt}(\mathbf{x}^{\intercal}\mathbf{x})\leq 0$ in view of the definition of $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)$. Therefore, all maximal solutions do not have finite escape times. Furthermore, according to (38), $\mathbf{x}^{+}=\mathbf{x},$ and from the definition of the update law in (34) and (37), it follows immediately that $\mathbf{J}(\mathcal{J})\subset\mathcal{K}$. Hence, the solutions to the hybrid closed-loop system (38) cannot leave $\mathcal{K}$ through jump and, as per [27, Proposition 6.10], all maximal solutions are complete.
-E Proof of Theorem 1
@@
Forward invariance and stability: The forward invariance of the obstacle-free set $\mathcal{K}$, for the hybrid closed-loop system (38), is immediate from Lemma 4. We next prove the stability of $\mathcal{A}$ using [27, Definition 7.1].